The sum of three consecutive intergers is 90.
a. Find the three numbers by letting x represent the first integer.
b. Find the three numbers by letting x represent the second number.

Answer by Alan3354(69443)

The sum of three consecutive intergers is 90.
a. Find the three numbers by letting x represent the first integer.
x + x+1 + x+2 = 90
-------------------
b. Find the three numbers by letting x represent the second number.
x-1 + x + x+1 = 90
-------------------
The best way:
90/3 = x, the middle number